Find the linear regression equation for the set of values. Round the slope and y-intercept to the nearest hundredth. {(5,-3.4), (10,-0.5), (15,2.1), (20,5)}
Please show the work so I can see how the problem was worked out. Thanks

Answer by rothauserc(4718)

regression equation is y = mx +b where m is slope and b is y intercept
m = (NΣXY - (ΣX)(ΣY)) / (NΣX^2 - (ΣX)^2)
b = (ΣY - m(ΣX)) / N
where N is number of paired values
(X, Y) are the individual points
Σ = summation
m = (4*109.5 -(50*3.2)) / (4*750 - 50^2) = 0.556 = 0.56
b = (3.2 - 0.56*50) / 4
b = -6.2
therefore our formula for the linear regression equation is
y = 0.56x -6.2